Dartmouth Distribution Warehouse makes deliveries of a large number of products to its customers. It is known that 80% of all the orders it receives from its customers are delivered on time. Let p^ be the proportion of orders in a random sample of 100 that are delivered on time. Find the probability that the value of p^ will be less than 0.84.

Round your answer to four decimal places.
Answer #1

Given,

p = 0.80 , n = 100

Using central limit theorem,

P(\hat{p} < p) = P( Z < \hat{p} - p / sqrt( p( 1 - p ) / n)

So,

P( \hat{p} < 0.84) = P( Z < 0.84 - 0.80 / sqrt( 0.80 * 0.20 / 100) )

= P(Z < 1)

= 0.8413